Frieda is keeping track of how many coupans she has savved on a piece of paper . she spilled some coffe on the paper. her lis now looks like this: blank represents where the coffee spill occured.
Started : 30 coupon After 1 week: after 2 weeks: after 3 weeks : 240 coupons
a. assume that the sequence of the number of coupons is arithmetic.How many coupon did she have after 1 week? 2 weeks?
b. Assume that the sequence of the number of coupons is geometric. How many coupons did she have after 1 week? after 2 weeks?

Answer by stanbon(75887)

Started : 30 coupon After 1 week: after 2 weeks: after 3 weeks : 240 coupons
a. assume that the sequence of the number of coupons is arithmetic.How many coupon did she have after 1 week? 2 weeks?
a(1) = 30 ; a(3) = 240
a(3) = a(1)+2d = 240
30 + 2d = 240
2d = 210
d = 105
After 1 week: 30
After 2 weeks: 30 + 105 = 135
-------------------------------------
b. Assume that the sequence of the number of coupons is geometric. How many coupons did she have after 1 week? after 2 weeks?
a(1) = 30: a(3) = 240
a(3) = a(1)*r^2 = 240
30r^2 = 240
r^2 = 8
r = 2sqrt(2)
---
after 1 week: 30
after 2 weeds: 20*2sqrt(2) = 40sqrt(2)
==========================================
Cheers,
Stan H.
=====================